# On the Utility of Accounting for Human Beliefs about AI Intention in Human-AI Collaboration

### v2

@@ 5. Conclusion
Additionally, our approach leverages historical data to construct models of human behavior and beliefs, which implicitly assumes that human behavior and beliefs remain stable over time. This may not hold true in real-world scenarios where human behavior can change in response to prolonged exposure to AI or due to evolving task contexts. Future research could focus on developing adaptive models that can learn and update human behavior and belief patterns in real-time to address this limitation. Investigating the scalability of these models to larger, more diverse datasets and environments is also important and interesting for assessing the generalizability and robustness of our approach.
Societal impacts. Our work underscores the importance of integrating human behavior and beliefs about AI into the design of AI systems to enhance human-AI collaboration. As AI systems become increasingly autonomous and capable, it is critical that these systems are designed with an understanding of human behavior and beliefs. This alignment can help improve trust, reduce errors in collaborative settings, and foster more effective interactions.
However, there are also potential ethical considerations and potential risks due to the dual-use of our approaches. In particular, the same methods used to improve human-AI collaboration could be misused, such as designing AI systems that manipulate or deceive users by exploiting their expectations or beliefs. This could undermine trust and create systems that harm user welfare. To mitigate these risks, one potential approach is to develop and enforce ethical guidelines that prioritize transparency in AI models. Providing clear and understandable explanations for AI behavior can help users build appropriate levels of trust and reliance. Moreover, promoting responsible AI design principles that incorporate user consent, control, and oversight are also important in ensuring that the technology serves societal benefits rather than causing harm.
{acks}
@@
Appendix A Additional Experiments
@@
A.1. Experiment Environment: Grid World with Multiple Goals
We use grid worlds of size of $6\times 6$ in both simulations and human experiments. Similar to the environment setup in Section 4.2, the grid world contains a start position, two goal positions, and some blocked positions that the player cannot enter. The player needs to move from the start position towards one of the goal positions. The player can choose to move {Up, Down, Right, Left}. The player will get a positive reward upon reaching the goal, and we set the maximum number of actions to be 20.
@@ A.1. Experiment Environment: Grid World with Multiple Goals
The interfaces of our experiments can be seen in Figure 6a, 6b, and 6c. The detailed descriptions of the experiment and interfaces are included in Appendix C.
@@
Figure 6. Human-subject experiment interfaces. In Experiment 4, each participant is asked to control the player to move to the goal (red star). In Experiment 5, each participant is provided a trace of the behavior by another agent, and is asked to infer which goal the agent is trying to reach. In Experiment 6, each participant is playing with an AI agent in separate environments. The participants only receive bonus rewards by reaching the same goal (red star or green triangle) as the AI agent.
A.2. Experiment 4: Evaluating Behavior Models in Single-Player MDP
@@
Table 3. The prediction accuracy for human behavior for different human models in Experiment 4.
@@
A.3. Experiment 5:
Evaluating Human Belief Models
@@ A.3. Experiment 5:
Table 4. Performance comparison between Bayesian inference framework assuming standard model and human behavior model.
@@
Figure 7. Human evaluations of belief inference accuracy regrading AI goals.
@@
A.4. Experiment 6:
Evaluating Collaborative AI Design
@@ A.4. Experiment 6:
Table 5. Simulation results of human-AI collaborative rewards in Experiment 6. Columns players are different AI agents, and row players are different simulated human models.
@@
Figure 8. Average collaborative reward in human-subject experiment 6.
A.5. Extensive Results of Behavioral Cloning
@@
Figure 9. Evaluation of model accuracy on human-subject experiment data through incorporation of action history.
Appendix B Implementation Details
@@
Selection of population size.
For FCP and MEP algorithms used in Section 4, the population size will directly influence the model performance. We vary the number of agents in our simulations as shown in Figure 10. The model trained with 8 partners is chosen for use in further simulations and human-subject experiments for comparative analysis.
Figure 10. Average team reward for models trained across different population sizes using the FCP and MEP algorithms.
Appendix C Human Experiment Details
@@
Figure 11. Human-subject experiment interfaces.
Table 6. Number of workers in each human-subject experiment.
@@
Table 7. Demographic information of all the participants in our experiments.

### v3

@@ 5 Conclusion and Discussion
Generalization and limitations. Our approach has proven effective through human-subject experiments in grid world environments of varying complexity. However, like most human-subject studies, the findings are limited to specific settings. To improve external validity, future work should evaluate the approach in more realistic domains, such as applications involving physical robots or complex multi-agent systems. Additionally, our approach relies on historical data to model human behavior and beliefs, assuming these remain stable over time. This may not hold true in real-world scenarios where human inference capability can change in response to prolonged exposure to AI or due to evolving task contexts. Future research could focus on developing adaptive models that can learn and update human behavior and belief patterns in real-time to address this limitation. Investigating the scalability of these models to larger, more diverse datasets and environments is also important and interesting for assessing the generalizability and robustness of our approach.
Acknowledgments
@@
Appendix A Related Work
Our work contributes to the growing body of research on human-AI collaboration [Carroll et al., 2019; Jaderberg et al., 2019; Bansal et al., 2021; Wilder et al., 2021; Myers et al., 2024]. For example, Bansal et al. [2021] demonstrate that optimizing AI in isolation may lead to suboptimal performance in human-AI collaboration, compared to training AI with consideration for human performance. Wilder et al. [2021] show that training AI systems to complement humans by performing better in areas where humans struggle results in improved collaborative performance. Carroll et al. [2019] illustrate that incorporating a human model, learned from human data, into the training of AI enhances performance when these AI systems collaborate with real humans. Our work extends this line of research by not only accounting for human behavior when designing AI but also incorporating human beliefs about AI intentions.
While the approaches in the above works often assume knowledge or data of humans, there have also been zero-shot coordination techniques being proposed to train cooperative AI agents when human data are not available. For example, Jaderberg et al. [2019] use population-based reinforcement learning to improve the robustness of trained AI agents. Strouse et al. [2021] develop a variety of self-play agents by varying random seeds of neural network initialization and various checkpoints, and train a best response model tailored to the agent population, discovering that humans prefer to partner with this model. Zhao et al. [2023] utilize the advances of maximum entropy reinforcement learning to encourage the diversity of agent populations and show the learned AI models could adapt to real humans. This line of work aims to design collaborative AI that is robust to diverse human behavior. However, they do not explicitly account for human beliefs over AI intentions during human-AI collaboration. In this work, we have included these approaches as baselines and show that accounting for human beliefs leads to better performance in our setting.
At a high level, our work aligns with the broader effort to understand and model humans in computational systems [Choudhury et al., 2019; Shah et al., 2019; Kwon et al., 2020; Chan et al., 2021; Reddy et al., 2021; Narayanan et al., 2022, 2023; Treiman et al., 2023, 2024, 2025]. More technically, for example, inverse reinforcement learning[Ng et al., 2000; Abbeel and Ng, 2004; Ramachandran and Amir, 2007] aims to infer reward functions in Markov decision processes (MDPs) by observing demonstrations of the optimal policy. When the demonstrator is human, these demonstrations may be noisy or exhibit behavioral biases. Several studies[Evans et al., 2016; Shah et al., 2019; Hughes et al., 2020; Zhi-Xuan et al., 2020] have addressed this by incorporating human behavioral biases into the inference process, aiming to infer both the rewards and biases simultaneously. Imitation Learning [Hussein et al., 2017] also seeks to develop models that can mimic human behavior based on demonstrations. Recently, there has been a growing amount of research that incorporates human models into computational and machine learning frameworks [Kleinberg and Oren, 2014; Tang and Ho, 2019, 2021; Masters et al., 2021a, b; Tang et al., 2021; Yu and Ho, 2022; Yu et al., 2023; Kasumba et al., 2024; Feng et al., 2024; Hong et al., 2024].
From the perspective of modeling human beliefs about others’ intentions, this has been discussed in the literature on level-$k$ reasoning [Stahl II and Wilson, 1994; Gill and Prowse, 2016] in economics and in the theory of mind [Premack and Woodruff, 1978; Chen et al., 2021; Westby and Riedl, 2023; Li et al., 2023]. For instance, Yu et al. [2006] found that human behavior aligns closely with level-$1$ and level-$2$ reasoning in cooperative games. Albrecht and Stone [2018] survey works on agents modeling the beliefs and intentions of other agents, distinguishing between methods for modeling stationary or changing agent behaviors. Bara et al. [2021] models human common ground in cooperative Minecraft tasks by leveraging theory of mind and develops AI models to infer human beliefs. Baker et al. [2011] propose a Bayesian theory of mind to describe how humans model joint beliefs of state in a partially observable MDP, conducting human experiments in the works of Baker et al. [2009, 2017]. Similarly, Wu et al. [2021] apply a human modeling approach in the Overcooked experiment, breaking down the task of delivering a dish into several subtasks, such as picking up an ingredient or placing an ingredient into a pot, where the player infers others’ intentions regarding these subtasks and selects their own accordingly before taking actions. Our work extends this line of research in two ways. First, our belief inference framework differs from previous theory of mind literature by modeling and incorporating human beliefs about AI intentions specifically, using a more realistic human model as the basis for inferring these intentions. Second, and more importantly, beyond modeling human beliefs over others’ intentions, our work focuses on the design of collaborative AI that account for human beliefs.
Appendix B Additional Experiments
@@
B.1 Experiment Environment: Grid World with Multiple Goals
We use grid worlds of size of $6\times 6$ in both simulations and human experiments. Similar to the environment setup in Section 2, the grid world contains a start position, two goal positions, and some blocked positions that the player cannot enter. The player needs to move from the start position towards one of the goal positions. The player can choose to move {Up, Down, Right, Left}. The player will get a positive reward upon reaching the goal, and we set the maximum number of actions to be 20.
@@ B.1 Experiment Environment: Grid World with Multiple Goals
The interfaces of our experiments can be seen in Figure 5a, 5b, and 5c. The detailed descriptions of the experiment and interfaces are included in Appendix D.
@@
Figure 5: Human-subject experiment interfaces. In Experiment 4, each participant is asked to control the player to move to the goal (red star). In Experiment 5, each participant is provided a trace of the behavior by another agent, and is asked to infer which goal the agent is trying to reach. In Experiment 6, each participant is playing with an AI agent in separate environments. The participants only receive bonus rewards by reaching the same goal (red star or green triangle) as the AI agent.
B.2 Experiment 4: Evaluating Behavior Models in Single-Player MDP
@@
Table 3: The prediction accuracy for human behavior for different human models in Experiment 4.
@@
B.3 Experiment 5: Evaluating Human Beliefs
@@
Figure 6: Human evaluations of belief inference accuracy regrading AI goals.
Table 4: Performance comparison between Bayesian inference framework assuming standard model and human behavior model.
@@
B.4 Experiment 6: Evaluating Collaborative AI
@@
Table 5: Simulation results of human-AI collaborative rewards in Experiment 6. Columns players are different AI agents, and row players are different simulated human models.
@@
Figure 7: Average collaborative reward in experiment 6.
B.5 Extensive Results of Behavioral Cloning
@@
Figure 8: Evaluation of model accuracy on human-subject experiment data through incorporation of action history.
In Experiment 1, we aggregate all data to train a behavioral cloning model. We also train models on human data collected with different AI teammates and evaluate them across treatments, observing consistent average accuracy ranging from $75.4\%$ to $78.9\%$. Since no significant differences were detected, we aggregate all the data to create a single human model. We also observed suboptimal performance from human participants in Experiment 1. Even when human participants are provided with suggested goals during data collection, there is only a $55.0\%$ chance that both players will reach the same type of goals across all treatments. Besides, there is a notable chance of two players colliding (about $15.8\%$) or ending in different types of goals (about $25.6\%$).
We also evaluate the model’s consistency accuracy in Experiment 3 and across different environment layouts. The data-driven model with a random goal achieves an accuracy of $61.15\%$, while incorporating belief inference improves accuracy to $73.19\%$ across all treatments, consistent with the trend shown in Figure 8. While the environment layout influences model accuracy, on average, the behavioral cloning model aligns with human behavior significantly better than assuming an optimal model. Figure 9 highlights scenarios where performance varies across treatments in Experiment 3. In the layout shown in Figure 9a, most AI types remain in their starting position and act only after the human player shows a clear preference for their goal. Here, the performance of Self-play AI is almost identical to Belief-AI, with no significant differences across other treatments. However, in the layout shown in Figure 9b, Belief-AI moves upward to signal its preference for the star, resulting in 10 out of 12 human players also reaching the star. In contrast, Self-play AI moves downward without clearly indicating its goal, leading to only 5 out of 13 participants aligning with the AI’s goal.
(a) Layout where Belief-AI (11 participants) and Self-play AI (9 participants) perform similarly.
(b) Layout where Belief-AI (12 participants) significantly outperforms Self-play AI (13 participants).
Figure 9: Examples of game layouts highlighting variations in performance across treatments.
Appendix C Implementation Details
@@
Figure 10: Average team reward for models trained across different population sizes using the FCP and MEP algorithms.
@@
Selection of population size. For FCP and MEP algorithms used in Section 4, the population size will directly influence the model performance. We vary the number of agents in our simulations as shown in Figure 10. The model trained with 8 partners is chosen for use in further simulations and human-subject experiments for comparative analysis.
Appendix D Human Experiment Details
@@
Figure 11: Human-subject experiment interfaces.
Table 6: Number of workers in each human-subject experiment.
@@
Table 7: Demographic information of all the participants in our experiments.